OpenStudy (anonymous):

Find the perimeter and area of a right triangle if one leg measures 8 cm and the other leg measures 15 cm. Students must show all work step by step to receive full credit for the problem. Include correct units with your solution.

OpenStudy (anonymous):

hypo = sqrt ( 64 + 225) = 17cm P = 17 + 8 + 15 = 40cm A = (1/2) 8 * 15 = 60 cm^2